# Learnable Tokenizer for LLM-based Generative Recommendation

### v1

@@ Abstract.
Harnessing Large Language Models (LLMs) for generative recommendation has garnered significant attention due to LLMs’ powerful capacities such as rich world knowledge and reasoning. However, a critical challenge lies in transforming recommendation data into the language space of LLMs through effective item tokenization. Existing approaches, such as ID identifiers, textual identifiers, and codebook-based identifiers, exhibit limitations in encoding semantic information, incorporating collaborative signals, or handling code assignment bias. To address these shortcomings, we propose LETTER (a LEarnable Tokenizer for generaTivE Recommendation), designed to meet the key criteria of identifiers by integrating hierarchical semantics, collaborative signals, and code assignment diversity. LETTER integrates Residual Quantized VAE for semantic regularization, a contrastive alignment loss for collaborative regularization, and a diversity loss to mitigate code assignment bias. We instantiate LETTER within two generative recommender models and introduce a ranking-guided generation loss to enhance their ranking ability. Extensive experiments across three datasets demonstrate the superiority of LETTER in item tokenization, thereby advancing the state-of-the-art in the field of generative recommendation.
@@ 1. Introduction
Harnessing Large Language Models (LLMs) for generative recommendation has emerged as a prominent avenue, drawing substantial research interest (Hua et al., 2023; Rajput et al., 2023). Leveraging the powerful abilities of LLMs (e.g., rich world knowledge, reasoning, and generalization), numerous efforts are investigating LLMs for generative recommendation, showcasing the potential of building foundational LLMs to move beyond natural language processing to information retrieval and recommendation domains (Geng et al., 2022; Cui et al., 2022). Technically, given user historically interacted items, generative recommendation aims to utilize LLMs to generate target items as recommendations. A key problem to encoding and generating items via LLMs is item tokenization, which indexes each item via an identifier (i.e., a token sequence), thereby bridging the gap between recommendation data and the language space of LLMs. As such, it is imperative to investigate item tokenization for LLM-based generative recommendation.
@@
Textual identifiers directly leverage item semantic information such as titles, attributes, or/and descriptions as item identifiers (Bao et al., 2023a; Hua et al., 2023). Nevertheless, such textual identifiers suffer from the following issues: 1) varying lengths of identifiers lead to bias, where longer identifiers are likely to have lower generation probabilities (see Figure 1 for empirical evidence); 2) semantic information is not hierarchically distributed from coarse to fine-grained in the token sequence of identifiers. This results in low generation probability if the first tokens of the target item are not closely aligned with user preference (e.g., “With” and “Be” in movie titles); 3) textual identifiers lack collaborative signals from user behaviors (Bao et al., 2023a). The token sequence of an identifier is solely determined by semantic information. As such, items with similar semantics yet differing collaborative signals exhibit similar token sequences, undermining the distinctiveness of items from the perspective of Collaborative Filtering (CF) (refer to Section 3.1).
Codebook-based methods adopt auto-encoders to encode item semantics into fixed-length hierarchical code sequences as identifiers (Rajput et al., 2023; Zheng et al., 2024). However, similar to textual identifiers, codebook-based identifiers still lack collaborative signals in code sequences (see Section 3.1 for details). Moreover, as shown in Figure 1, the assignment of codes to items is naturally imbalanced, leading to the item generation bias where items with high-frequency codes are more readily generated.
@@
Integrating hierarchical semantics into fixed-length identifiers, where the token sequence initially encodes broad and coarse-grained semantics, progressively transitioning into more refined, fine-grained details. This aligns with the autoregressive generation characteristics of generative recommendation.
@@
To this end, we propose a LEarnable Tokenizer for generaTivE Recommendation named LETTER. LETTER incorporates three kinds of regularization to enhance the codebook-based identifiers. In particular, semantic regularization first integrates Residual Quantized VAE (RQ-VAE) (Lee et al., 2022) to translate the item semantic information into the fixed-length hierarchical identifiers (Rajput et al., 2023). Built on this, collaborative regularization utilizes a contrastive alignment loss to align semantic quantized embeddings in RQ-VAE with CF embeddings from a well-trained CF model (e.g., LightGCN (He et al., 2020)). Moreover, LETTER introduces a diversity loss to enhance code embedding diversity to alleviate code assignment bias and item generation bias. We instantiate LETTER on two representative generative recommender models and propose a ranking-guided generation loss to boost the ranking ability of the generative models theoretically. Extensive experiments on three datasets with in-depth investigation have validated that LETTER can achieve superior item tokenization by simultaneously considering hierarchical semantics, collaborative signals, and code assignment diversity in identifiers.
@@
We comprehensively analyze the necessary features of an ideal identifier. Accordingly, we propose a novel learnable tokenizer named LETTER to adaptively learn identifiers encompassing hierarchical semantics, collaborative signals, and code assignment diversity.
@@ 2. Item Tokenization
Identifier length bias. Textual identifiers using unequal-length titles and descriptions face the bias issue that longer identifiers unfairly have lower generation probabilities. We employ BIGRec (Bao et al., 2023a), a representative model with textual identifiers, to conduct experiments in the Instruments dataset (Ni et al., 2019) for empirical verification: we compare items with different identifier lengths and match items’ occurrence frequency in the training data to avoid the influence of item popularity111Popular item identifiers in the training data are more likely to be generated (Jiang et al., 2024).. By comparing the generation frequency in Figure 1(a), we find that a longer identifier (length=10) has a lower generation probability given the same popularity, validating the existence of identifier length bias.
Non-hierarchical semantics. Semantic information is not hierarchically distributed from coarse to fine-grained in textual identifiers. Since LLMs generate identifiers in an autoregressive manner, the initial tokens in identifiers are expected to represent the coarse-grained and pivotal semantics.
@@
Code assignment bias. codebook-based methods additionally suffer from imbalanced code assignment, leading to item generation bias. We utilize TIGER (Rajput et al., 2023) in Instruments for empirical investigation. Specifically, we split item identifier codes into five groups based on their assignment popularity in the training data and obtain the generation frequency of TIGER. From Figure 1(b), we can observe that popular item identifiers are more likely to be generated, amplifying the generation bias.
In light of these, we consider three essential objectives to pursue an ideal identifier: 1) hierarchical semantic integration with fixed-length token sequence, 2) collaborative signals integration in the token sequence, and 3) high diversity of the code assignments.
@@ 3.1. LETTER
To achieve the three objectives for item tokenization, we build LETTER for LLM-based generative recommendation. In particular, as depicted in Figure 3, LETTER can generate fixed-length identifiers, i.e., code sequences, with hierarchical semantics. Moreover, we introduce collaborative regularization to endow the code sequence with collaborative signals, and diversity regularization to alleviate the code assignment bias issue during tokenizer training.
Semantic regularization. To pursue identifiers with hierarchical semantics with a fixed length, we develop LETTER based on RQ-VAE (Lee et al., 2022), which encodes the hierarchical item semantics into a code sequence. As a multi-level embedding quantizer, RQ-VAE recursively quantizes semantic residuals using a fixed number of codebooks, thus naturally achieving the fixed length identifier with coarse to fine-grained semantics. As shown in Figure 3, the learnable tokenizer generates the hierarchical semantic identifier via two steps: 1) semantic embedding extraction, and 2) semantic embedding quantization.
Semantic embedding extraction. Given an item with its content information such as titles and descriptions, we first extract the semantic embedding through a pre-trained semantic extractor, e.g., LLaMA2-7B (Touvron et al., 2023). The semantic embedding is then compressed into the latent semantic embedding through an encoder:
@@
where is the stop-gradient operation (Van Den Oord et al., 2017), and is the coefficient to balance the strength between the optimization of code embeddings and encoder. Specifically, the reconstruction loss aims to maintain the essential semantics information in the latent space such that the semantic embedding can be reconstructed from the quantized embedding. Besides, aims to reduce the residual error for all levels and jointly train the encoder, and code embeddings (Van Den Oord et al., 2017). By semantic regularization, the code sequence encodes hierarchical semantics, facilitating next-item generation and cold-start generalization.
Collaborative regularization. To inject collaborative signals into the code sequence instead of only code embeddings, we introduce collaborative regularization to align the quantized embedding and the CF embedding via contrastive learning. Specifically, we utilize a well-trained CF recommender model (e.g., SASRec (Kang and McAuley, 2018) and LightGCN (He et al., 2020)) to obtain CF embeddings of items, which is then aligned with the quantized embedding through a CF loss. The CF loss is formulated as
@@
Diversity regularization. To address the code generation bias, a promising solution is to learn a more uniform distribution of the code embeddings. As illustrated in Figure 5(a), a biased code embedding distribution is more likely to cause biased code assignment in the latent space, where some codes are assigned to more items. In contrast, a uniform code embedding distribution facilitates more balanced code assignments (Figure 5(b)). Based on the above insights, we aim to improve the code embedding diversity for diverse code assignments.
@@ 3.2.1. Training
where is the adjustable temperature to penalize the hard-negative samples, refers to the -th token of , represents the token sequence preceding , denotes the learnable parameters of the generative model, and is the token vocabulary222The token vocabulary is extended by adding the codes from LETTER..
@@ 4.1.1. Datasets
We use three real-world recommendation datasets from different domains: 1) Instruments is from the Amazon review datasets (Ni et al., 2019), which contains user interactions with rich music gears. 2) Beauty contains user interactions with extensive beauty products from Amazon review datasets (Ni et al., 2019). 3) Yelp333https://www.yelp.com/dataset. is a popular dataset comprising business interactions on the Yelp platform. We adopt the pre-processing techniques from previous work (Kang and McAuley, 2018), discarding sparse users and items with interactions less than 5. We consider sequential recommendation setting444We focus on sequential recommendation setting due to its high practicality and significant potential in real-world applications. and use leave-one-out strategy (Rajput et al., 2023; Zheng et al., 2024) to split datasets. For training, we follow (Hua et al., 2023; Zheng et al., 2024) to restrict the number of items in a user’s history to 20.
@@ 4.1.2. Baselines
We compare LETTER with competitive baselines within two groups of work: traditional recommender models and LLM-based generative recommender models with different types of identifiers (i.e., ID, textual, and codebook-based identifiers): 1) MF (Rendle et al., 2012) decomposes the user-item interactions into the user embeddings and the item embeddings in the latent space. 2) Caser (Tang and Wang, 2018) employs convolutional nerual networks to capture user’s spatial and positional information. 3) HGN (Ma et al., 2019) utilizes graph neural networks to learn user and item representations for predicting user-item interactions. 4) BERT4Rec (Sun et al., 2019) leverages BERT’s pre-trained language representations to capture semantic user-item relationships. 5) LightGCN (He et al., 2020) is a lightweight graph convolutional network model focusing on high-order connections between users and items. 6) SASRec (Kang and McAuley, 2018) employs self-attention mechanisms to capture long-term dependencies in user interaction history. 7) BIGRec (Bao et al., 2023a) is an LLM-based generative recommender model using textual identifiers, with each item represented by its title. 8) P5-SemID (Hua et al., 2023) leverages item metadata (e.g., attributes) to construct ID identifiers for LLM-based generative recommender model. 9) P5-CID (Hua et al., 2023) incorporates collaborative signals into the identifier for LLM-based generative recommender models through a spectral clustering tree derived from item co-appearance graphs. 10) TIGER (Rajput et al., 2023) introduces codebook-based identifiers via RQ-VAE, which quantizes semantic information into code sequence for LLM-based generative recommendation. 11) LC-Rec (Zheng et al., 2024) uses codebook-based identifiers and auxiliary alignment tasks to better utilize knowledge in LLMs by connecting generated code sequences with natural language.
4.1.3. Evaluation Settings
We use two metrics: top- Recall (R@) and NDCG (N@) with = 5 and 10, following (Rajput et al., 2023).
@@ 4.1.4. Implementation Details
We instantiate LETTER on two representative LLM-based generative recommender models, i.e., TIGER (Rajput et al., 2023) and LC-Rec (Zheng et al., 2024). As for TIGER, since the official implementations have not been released, we follow the paper for implementation. For LC-Rec, we perform the parameter-efficient fine-tuning technique LoRA (Hu et al., 2021) to fine-tune LLaMA2-7B (Touvron et al., 2023). To obtain item semantic embedding, we follow (Zheng et al., 2024) to adopt LLaMA2-7B to encode item content information. We obtain 32-dimensional item embedding from SASRec (Kang and McAuley, 2018) for collaborative regularization and set cluster number at 10 for diversity regularization. All the experiments are conducted on 4 NVIDIA RTX A5000 GPUs.
@@ 4.2. Overall Performance (RQ1)
We evaluate LETTER on two SOTA LLM-based generative recommender models on three real-world datasets. The performances of baselines and LETTER instantiated on TIGER (LETTER-TIGER) and LC-REC (LETTER-LC-Rec) are reported in Table 1, from which we have the following observations:
Among the LLM-based models with ID identifiers (P5-CID and P5-SemID), P5-CID usually yields better performance than P5-SemID over the three datasets. This is because P5-SemID assigns item identifiers based on the item categories, e.g., string and keyboard for instrument products, which might 1) fail to capture fine-grained semantic information, and 2) hinder the learning of collaborative behavior in LLMs’ fine-tuning stage due to the misalignment between semantic and collaborative signals (cf. Figure 4 in Section 3.1). In contrast, P5-CID leverages CF information from the item co-appearance graph to assign identifiers, which is beneficial for LLM-based recommender models to capture collaborative patterns in user behaviors.
Among all LLM-based baselines with different kinds of identifiers, models with codebook-based identifiers (TIGER and LC-Rec) outperform models with ID identifiers (P5-SemID and P5-CID) and that with textual identifiers (BIGRec) in most cases. This is attributed to the incorporation of hierarchical semantics with different granularities via RQ-VAE, which distinguishes between similar items more effectively by grasping fine-grained details. Meanwhile, the inferior performance of BIGRec with textual identifiers is probably due to 1) the item generation bias caused by identifier length bias, and 2) the potential misalignment between items’ similar semantics and dissimilar interactions, thus hurting the learning of collaborative signals for accurate recommendation.
LETTER consistently exhibits significant improvements for the backend models (TIGER and LC-Rec) across three datasets, validating the effectiveness of our method. The superior performance is attributed to: 1) the CF integration during code assignment, which aligns the collaborative signals and the semantic embeddings, thereby encouraging items with similar interactions or semantics to have similar code sequence and addressing the misalignment issue between semantics and CF information (refer to Section 4.3.4 for further analysis of CF integration). 2) The improved diversity of token assignments, which is achieved by regularizing the representation space of code embeddings, thereby alleviating item generation bias caused by code assignment bias (refer to Section 4.3.2 for detailed analysis).

### v2

# Learnable Item Tokenization for Generative Recommendation

@@ Abstract.
Utilizing powerful Large Language Models (LLMs) for generative recommendation has attracted much attention. Nevertheless, a crucial challenge is transforming recommendation data into the language space of LLMs through effective item tokenization. Current approaches, such as ID, textual, and codebook-based identifiers, exhibit shortcomings in encoding semantic information, incorporating collaborative signals, or handling code assignment bias. To address these limitations, we propose LETTER (a LEarnable Tokenizer for generaTivE Recommendation), which integrates hierarchical semantics, collaborative signals, and code assignment diversity to satisfy the essential requirements of identifiers. LETTER incorporates Residual Quantized VAE for semantic regularization, a contrastive alignment loss for collaborative regularization, and a diversity loss to mitigate code assignment bias. We instantiate LETTER on two models and propose a ranking-guided generation loss to augment their ranking ability theoretically. Experiments on three datasets validate the superiority of LETTER, advancing the state-of-the-art in the field of LLM-based generative recommendation.
@@ 1. Introduction
Harnessing Large Language Models (LLMs) for generative recommendation has emerged as a prominent avenue, drawing substantial research interest (Hua et al., 2023; Rajput et al., 2023). Leveraging the powerful abilities of LLMs (e.g., rich world knowledge, reasoning, and generalization), numerous efforts are investigating LLMs for generative recommendation, showcasing the potential of building foundational LLMs to move beyond natural language processing to information retrieval and recommendation domains (Geng et al., 2022; Cui et al., 2022). As illustrated in Figure 1, given user historically interacted items, generative recommendation aims to utilize LLMs to generate target items as recommendations. A key problem to encoding and generating items via LLMs is item tokenization, which indexes each item via an identifier (i.e., a token sequence), thereby bridging the gap between recommendation data and the language space of LLMs. As such, it is imperative to investigate item tokenization for LLM-based generative recommendation.
@@
Textual identifiers directly leverage item semantic information such as titles, attributes, or/and descriptions as item identifiers (Bao et al., 2023a; Hua et al., 2023). Nevertheless, such textual identifiers suffer from the following issues: 1) semantic information is not hierarchically distributed from coarse to fine-grained in token sequence of the identifiers. This results in low generation probability if the first tokens of the target item are not closely aligned with user preference (e.g., “With” and “Be” in movie titles); 2) textual identifiers lack collaborative signals from user behaviors (Bao et al., 2023a). The token sequence of an identifier is solely determined by semantic information. As such, items with similar semantics yet differing collaborative signals exhibit similar token sequences, undermining the distinctiveness of items from the perspective of Collaborative Filtering (CF) (refer to Section 2 and Figure 2).
Codebook-based methods adopt auto-encoders to encode item semantics into hierarchical code sequences as identifiers (Rajput et al., 2023; Zheng et al., 2024). However, similar to textual identifiers, codebook-based identifiers still lack collaborative signals in code sequences (see Section 2). Moreover, as shown in Figure 3, the assignment of codes to items is naturally imbalanced, causing the item generation bias where items with high-frequency codes are more readily generated.
@@
Integrating hierarchical semantics into identifiers, where the token sequence initially encodes broad and coarse-grained semantics, progressively transitioning into more refined, fine-grained details. This aligns with the autoregressive generation characteristics of generative recommendation.
@@
To this end, we propose a LEarnable Tokenizer for generaTivE Recommendation named LETTER. LETTER incorporates three kinds of regularization to enhance the codebook-based identifiers. In particular, semantic regularization first integrates Residual Quantized VAE (RQ-VAE) (Lee et al., 2022) to translate the item semantic information into the hierarchical identifiers (Rajput et al., 2023). Built on this, collaborative regularization utilizes a contrastive alignment loss to align semantic quantized embeddings in RQ-VAE with CF embeddings from a well-trained CF model (e.g., LightGCN (He et al., 2020)). Moreover, LETTER introduces a diversity loss to enhance code embedding diversity to alleviate code assignment bias and item generation bias. We instantiate LETTER on two representative generative recommender models and propose a ranking-guided generation loss to boost the ranking ability of these generative models theoretically. Extensive experiments on three datasets with in-depth investigation have validated that LETTER can achieve superior item tokenization by simultaneously considering hierarchical semantics, collaborative signals, and code assignment diversity in identifiers. For reproducibility, we release our code and data at https://github.com/HonghuiBao2000/LETTER.
@@
We comprehensively analyze the necessary features of an ideal identifier. Accordingly, we propose a novel learnable tokenizer named LETTER, aimed at adaptively learning identifiers that encompass hierarchical semantics, collaborative signals, and code assignment diversity.
@@ 2. Item Tokenization
Non-hierarchical semantics. As mentioned in Section 1, semantic information is not hierarchically distributed from coarse to fine-grained in textual identifiers.
@@
Code assignment bias. Codebook-based methods additionally suffer from imbalanced code assignment, leading to item generation bias. We utilize TIGER (Rajput et al., 2023) on Instruments dataset for empirical investigation. Specifically, we split item identifier codes into five groups based on their assignment popularity in the training data and obtain the generation frequency of TIGER. From Figure 3, we can observe that popular item identifiers are more likely to be generated, amplifying the generation bias.
In light of these, we consider three essential objectives to pursue an ideal identifier: 1) hierarchical semantic integration with the token sequence, 2) collaborative signals integration in the token sequence, and 3) high diversity of the code assignment.
@@ 3.1. LETTER
To achieve the three objectives for item tokenization, we build LETTER for LLM-based generative recommendation. In particular, as depicted in Figure 4, LETTER can generate identifiers, i.e., code sequences, with hierarchical semantics. Moreover, we introduce collaborative regularization to endow the code sequence with collaborative signals, and diversity regularization to alleviate the code assignment bias issue during tokenizer training.
Semantic regularization. To pursue identifiers with hierarchical semantics, we develop LETTER based on RQ-VAE (Lee et al., 2022), which encodes the hierarchical item semantics into a code sequence. As a multi-level embedding quantizer, RQ-VAE recursively quantizes semantic residuals using a fixed number of codebooks, thus naturally achieving identifier with coarse to fine-grained semantics. As shown in Figure 4, the learnable tokenizer generates the hierarchical semantic identifier via two steps: 1) semantic embedding extraction, and 2) semantic embedding quantization.
Semantic embedding extraction. Given an item with its content information such as titles and descriptions, we first extract the semantic embedding through a pre-trained semantic extractor, e.g., LLaMA-7B (Touvron et al., 2023). The semantic embedding is then compressed into the latent semantic embedding through an encoder .
@@
where is the stop-gradient operation (Van Den Oord et al., 2017), and is the coefficient to balance the strength between the optimization of code embeddings and encoder. Specifically, the reconstruction loss aims to maintain the essential semantics information in the latent space such that the semantic embedding can be reconstructed from the quantized embedding. Besides, reduces the residual error for all levels and jointly trains the encoder, and code embeddings (Van Den Oord et al., 2017). By semantic regularization, the code sequence encodes hierarchical semantics, facilitating coarse-grained to fine-grained generation and cold-start generalization.
Collaborative regularization. To inject collaborative signals into the code sequence instead of only code embeddings, we introduce collaborative regularization to align the quantized embedding and the CF embedding via contrastive learning. Specifically, we utilize a well-trained CF recommender model (e.g., SASRec (Kang and McAuley, 2018) and LightGCN (He et al., 2020)) to obtain CF embeddings of items, which is then aligned with the quantized embedding through a CF loss:
@@
Diversity regularization. To address the code generation bias, it is promising to learn a more uniform distribution of the code embeddings. As illustrated in Figure 5(a), a biased code embedding distribution is more likely to cause biased code assignment in the latent space, where some codes are assigned to more items. In contrast, a uniform code embedding distribution facilitates more balanced code assignment (Figure 5(b)). In light of these, we aim to improve the code embedding diversity for diverse code assignments.
@@ 3.2.1. Training
where is the adjustable temperature to penalize the hard-negative samples, refers to the -th token of , represents the token sequence preceding , denotes the learnable parameters of the generative model, and is the token vocabulary111The token vocabulary is extended by adding the codes from LETTER..
@@ 4.1.1. Datasets
We use three real-world recommendation datasets from different domains: 1) Instruments is from the Amazon review datasets (Ni et al., 2019), which contains user interactions with rich music gears. 2) Beauty contains user interactions with extensive beauty products from Amazon review datasets (Ni et al., 2019). 3) Yelp222https://www.yelp.com/dataset. is a popular dataset comprising business interactions on the Yelp platform. We adopt the pre-processing techniques from previous work (Kang and McAuley, 2018; Rajput et al., 2023), discarding sparse users and items with interactions less than 5. We consider sequential recommendation setting333We focus on sequential recommendation setting due to its high practicality and significant potential in real-world applications. and use leave-one-out strategy (Rajput et al., 2023; Zheng et al., 2024) to split datasets. For training, we follow (Hua et al., 2023; Zheng et al., 2024) to restrict the number of items in a user’s history to 20.
@@ 4.1.2. Baselines
We compare LETTER with competitive baselines within two groups of work, traditional recommender models and LLM-based generative recommender models, with different types of identifiers (i.e., ID, textual, and codebook-based identifiers): 1) MF (Rendle et al., 2012) decomposes the user-item interactions into the user embeddings and the item embeddings in the latent space. 2) Caser (Tang and Wang, 2018) employs convolutional nerual networks to capture user’s spatial and positional information. 3) HGN (Ma et al., 2019) utilizes graph neural networks to learn user and item representations for predicting user-item interactions. 4) BERT4Rec (Sun et al., 2019) leverages BERT’s pre-trained language representations to capture semantic user-item relationships. 5) LightGCN (He et al., 2020) is a lightweight graph convolutional network model focusing on high-order connections between users and items. 6) SASRec (Kang and McAuley, 2018) employs self-attention mechanisms to capture long-term dependencies in user interaction history. 7) BIGRec (Bao et al., 2023a) is an LLM-based generative recommender model using textual identifiers, with each item represented by its title. 8) P5-TID (Hua et al., 2023) uses items’ title as textual identifiers for LLM-based generative recommender model. 9) P5-SemID (Hua et al., 2023) leverages item metadata (e.g., attributes) to construct ID identifiers for LLM-based generative recommender model. 10) P5-CID (Hua et al., 2023) incorporates collaborative signals into the identifier for LLM-based generative recommender models through a spectral clustering tree derived from item co-appearance graphs. 11) TIGER (Rajput et al., 2023) introduces codebook-based identifiers via RQ-VAE, which quantizes semantic information into code sequence for LLM-based generative recommendation. 12) LC-Rec (Zheng et al., 2024) uses codebook-based identifiers and auxiliary alignment tasks to better utilize knowledge in LLMs by connecting generated code sequences with natural language.
Evaluation settings. We use two metrics: top- Recall (R@) and NDCG (N@) with = 5 and 10, following (Rajput et al., 2023).
@@ 4.1.3. Implementation Details
We instantiate LETTER on two representative LLM-based generative recommender models, i.e., TIGER (Rajput et al., 2023) and LC-Rec (Zheng et al., 2024). As for TIGER, since the official implementations have not been released, we follow the paper for implementation. For LC-Rec, we perform the parameter-efficient fine-tuning technique LoRA (Hu et al., 2021) to fine-tune LLaMA-7B (Touvron et al., 2023). To obtain item semantic embedding, we follow (Zheng et al., 2024) to adopt LLaMA-7B for encoding item content information. We obtain 32-dimensional item embedding from SASRec (Kang and McAuley, 2018) for collaborative regularization and set cluster number at 10 for diversity regularization. All the experiments are conducted on 4 NVIDIA RTX A5000 GPUs.
@@ 4.2. Overall Performance (RQ1)
We evaluate LETTER on two SOTA LLM-based generative recommender models. The performances of LETTER instantiated on TIGER (LETTER-TIGER) and LC-REC (LETTER-LC-Rec) are reported in Table 1, from which we have the following observations:
Among the LLM-based models with ID identifiers (P5-CID and P5-SemID), P5-CID usually yields better performance than P5-SemID over the three datasets. This is because P5-SemID assigns item identifiers based on the item categories, e.g., “string” and “keyboard” for instrument products, which might 1) fail to capture fine-grained semantic information and 2) hinder the learning of collaborative behavior in LLMs’ fine-tuning stage due to the misalignment between semantic and collaborative signals (cf. Figure 2 in Section 2). In contrast, P5-CID leverages collaborative signals from the item co-appearance graph to assign identifiers, which is beneficial for LLM-based recommender models to capture collaborative patterns in user behaviors.
Among all LLM-based baselines with different kinds of identifiers, models with codebook-based identifiers (TIGER and LC-Rec) outperform models with ID identifiers (P5-SemID and P5-CID) and that with textual identifiers (BIGRec and P5-TID) in most cases. This is attributed to the incorporation of hierarchical semantics with different granularities via RQ-VAE, which distinguishes between similar items more effectively by grasping fine-grained details. Meanwhile, the inferior performance of BIGRec and P5-TID with textual identifiers is probably due to the potential misalignment between items’ similar semantics and dissimilar interactions, thus hurting the learning of collaborative signals for accurate recommendation.
LETTER consistently exhibits significant improvements for the backend models (TIGER and LC-Rec) across three datasets, validating the effectiveness of our method. The superior performance is attributed to: 1) the CF integration during code assignment, which aligns the collaborative signals and the semantic embeddings, thereby encouraging items with similar interactions or semantics to have similar code sequence and addressing the misalignment issue between semantics and collaborative signals (refer to Section 4.3.4 for further analysis of CF integration). 2) The improved diversity of token assignments, which is achieved by regularizing the representation space of code embeddings, thereby alleviating item generation bias caused by code assignment bias (refer to Section 4.3.2 for detailed analysis).